A 71 kg ball (Ball A) traveling to the right at 7 m/s collides with a 22 kg ball (Ball B) at rest. After the collision, the 22 kg ball (Ball B) is traveling to the right at 44 m/s. What is the final velocity of Ball A?

(please show as much work as possible...thank you!)
Answer #1

Using the conservation of momentum

M1U1 = M1V1 + M2V2

71*7 = 22*44 + 71*V2

V2 = -471/71

= - 6.633 m/s ( - sign indicates moving to left)

Therefore final velocity of A = -6.633 m/s